# Severity Controlled Text-to-Image Generative Model Bias Manipulation

### v1

@@ 4.1 Results
Across all tasks, our goal is to manipulate an output image, shifting the bias from class $\mathbb{A}$ to $\mathbb{B}$ by utilising $\mathcal{S}$ variables to control manipulation severity. Our precise prompt engineering results in Table I show that the relationships between $\mathcal{S}$ and our evaluation metrics are consistent for all four common class combinations, evidencing that we have achieved fine control over the language model embedding output. We also see that extrapolating the embedding beyond $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ is effective as visualized in Fig. 5,6 and supported quantitatively in Table I. For $\mathcal{S}<0$ and $\mathcal{S}>1$, we can clearly see that the embedding is still within clusters $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ respectively. Where $\mathcal{P}_{\mathbb{A}}\cong\mathcal{P}_{\mathbb{B}}\cong 0.5$, this implies that applying the relevant transformation with the corresponding $\mathcal{S}_{i}$ has shifted $\mathbf{X}$ to the border of clusters $\mathbb{A}$ and $\mathbb{B}$ i.e. a precisely-engineered, empirically-derived, “half-dog, half-cat”.
For our gender, age and race embedding manipulation, we report the quantitative results in Table II and present qualitative examples in Fig. 7. We observe that the effectiveness of prompt extrapolation (where $\mathcal{S}<0.0~{}\cap~{}\mathcal{S}>1.0$) varies depending on the task, particularly in some cases where $\mathcal{S}=-3.0,3.0$. We find that examples of extreme extrapolation sometimes generated illegible images for age and race-related embedding transformations as shown in Fig. 7 (b). We also observe that our method is robust to long, complex input prompts, as evidenced in Fig. 7 (b), where the appropriate feature manipulations in the image were apparent irrespective of the length of the input prompt.
@@
Through results of the $N_{20}$ and $N_{100}$ social bias tuning experiments in Table III, we immediately observe that $\mathcal{S}_{i}$ values are dependent on the amount of seeds used for image generation. Covering a broader range of samples improves generalizability but at the cost of computation time. We opt for using labels to describe $\mathcal{S}_{i}$ and $\mathcal{P}_{i}$ for clarity in Table III. However this is just a matter of syntax and $\mathcal{S}_{male}\rightarrow\mathcal{S}_{asian}$ are identical to $\mathcal{S}_{1}\rightarrow\mathcal{S}_{7}$ derived previously.
@@
TABLE III: Social bias mitigation results. For $N_{20}$ and $N_{100}$ random seed experiments, we tune $\mathcal{S}_{i}$ to balance gender, age and racial representations, where $\mathcal{S}_{i}$ represents how much the embedding has been shifted toward the $i^{th}$ centroid based on Eq. (5). We exploit a CLIP VC model to report the frequency/probability ‘$\mathcal{P}_{i}$’ of generated images being classified as class $i$. Fig. 4 presented a visual representation of how we optimize $\mathcal{S}_{i}$.
Task Dog - Cat Bird - Plane
$\mathbf{x}_{T}$ Photo Picture Image View - Photo Picture Image View
$\mathcal{S}$ 0 0.5 0.75 1 1.25 0 0.5 0.75 1 1.25
ASRVC 0 0.049 0.848 1 1 0.005 0.140 0.929 1 0.989
ASRVL 0.011 0.115 0.860 0.987 0.982 0.043 0.127 0.626 0.400 0.495
$\mathcal{P}_{\mathbb{B}}$ 0.000 0.059 0.837 0.988 0.993 0.005 0.141 0.921 0.999 0.980
Task Horse - Deer Car - Truck
ASRVC 0 0 0.249 0.9 1 0 0.1 0.749 1 0.999
ASRVL 0 0.001 0.006 0.400 0.863 0.024 0.186 0.599 1 0.978
$\mathcal{P}_{\mathbb{B}}$ 0.000 0.003 0.244 0.916 0.996 0.001 0.119 0.731 0.980 0.993
@@
TABLE IV: Semantically-null backdoor attack results. For the four tasks, we define semantically-null trigger clusters $\mathbb{A}_{i}$, each hosting a unique severity value $\mathcal{S}_{i}$. Upon detection of a trigger ‘$x_{T_{i}}$’ in the prompt, we apply the shift to the embedding output based on Eq. (7). We report VL and VC attack success rate (ASRVL and ASRVC, respectively) and classifier confidence/probability for class $\mathbb{B}$ $\mathcal{P}_{\mathbb{B}}$, where $\mathcal{P}_{\mathbb{A}}=1-\mathcal{P}_{\mathbb{B}}$.
Figure 8: Qualitative backdoor attack results, using the same random seed (de-noising conditions) to generate images for each task. We prepend $x_{T_{i}}$ to the base input prompt in our experiments e.g. “a photo of…”. (a) prompt class = dog, $\mathbb{B}$ = cat. (b) prompt class = bird, $\mathbb{B}$ = plane.
Figure 9: Qualitative backdoor attack results, using the same random seed (de-noising conditions) to generate images for each task. We prepend $x_{T_{i}}$ to the base input prompt in our experiments e.g. “a photo of…”. (a) prompt class = horse, $\mathbb{B}$ = deer. (b) prompt class = car, $\mathbb{B}$ = truck.
In Table III, we report the post-tuning results for each social marker (gender, age, race) and compare it to the base model, where $\mathcal{S}_{1\rightarrow 7}=0$. For the $N_{100}$ case, we see that gender biases were evenly distributed in the base model, without making any changes to $\mathcal{S}_{male}$, $\mathcal{S}_{female}$. Whereas for the $N_{20}$ case, there is a clear gender imbalance that required tuning. As predicted, changing $\mathcal{S}_{i}$ variables will have an impact on class representations at each tuning stage, leading to the final step where the biases are more evenly distribution than when compared to the base models. Clemmer et al. reported less than 2.5% gender bias and less than 10% ethnicity biases in T2I models [4]. Similarly, Feng et al. reported a convergence of around 50% gender probability distribution for binary gender classes (less than 7% bias) in [6]. Comparing our results in Table III, we see that our mitigation method is comparable to the state of the art works as we report less than 10% gender, age and racial bias.
@@
Given the attack nature of backdoors, ASR metrics are important to demonstrate that our approach is effective. Logically, we hypothesized that ASRVL, ASRVC and $\mathcal{P}_{\mathbb{B}}$ are all proportional to $\mathcal{S}$ in the range $0\leq\mathcal{S}\leq 1$. In comparison, we expect that $\mathcal{P}_{\mathbb{A}}$ is inversely proportional to $\mathcal{S}$ in the range $0\leq\mathcal{S}\leq 1$. In Table IV, we observe that our hypotheses were correct for all four CIFAR-10 class pairs. Through Fig. 8 and 9 as expected, the output images converge on their respective $c_{\mathbb{B}}$ as $\mathcal{S}\rightarrow 1$. The weakest performing task was the Horse-Deer attack, where we observe that unlike the others, it is the only case where ASRVC is not 100% at $\mathcal{S}=1$ This may be due to the similarities between the two labelled points. For others, we see that the attack is quite effective even at $\mathcal{S}=0.75$. Overall, we demonstrate an effective, computationally-efficient backdoor injection method, where we manipulate $\mathbf{X}$ and have control over the severity through a semantically-null trigger space $\mathbb{A}_{i}$ and the corresponding severity - $\mathcal{S}_{i}$.
@@
For comparison, in [7], Vice et al. reported a backdoor ASR range of 47.2% to 87.9%. Chen et al. achieved a 79.3% to 99.6% ASR in [14] when targeting the diffusion/generative model component. Struppek et al. targeted embedded language models and reported an attack accuracy of around 90% [15]. Zhai et al. reported a 60.1% to 98.8% ASR range when conducting their backdoor implementations [30]. Thus, not only can we effectively control the severity of our attacks in real-time, achieving mean ASRVC and ASRVL scores of 97.5% and 69.2% respectively proves that at $\mathcal{S}=1.0$, our attacks are comparable to state-of-the-art works. Furthermore, by extrapolating $\mathbf{X}$ beyond $c_{\mathbb{B}}$ at $\mathcal{S}=1.25$, our reported ASRVC is higher than other methods [14, 15, 7, 30].
@@ 5 Conclusion
We have presented a computationally efficient, language embedding interpolation and extrapolation method which leverages high-dimensionality clusters and vector algebra to shift language model outputs and ultimately, control bias characteristics of downstream tasks. We target text-to-image models and show that bias manipulation is possible without fine-tuning or gaining access to network weights and parameters. Our approach is multi-faceted and we highlight applications for: (i) precise prompt engineering, (ii) improving social representations and, (iii) using semantically-null triggers to inject malicious backdoors. Our method provides more precise control of prompts in the embedding space and is comparable to state-of-the-art bias mitigation and backdoor attack methods. Through interpolation and extrapolation, manipulated embeddings allow us to generate images that would be near-impossible to describe using textual inputs.
@@ 6 Acknowledgements
This research was supported by National Intelligence and Security Discovery Research Grants (project #NS220100007), funded by the Department of Defence, Australia.

### v2

@@ 4 Experiments
Results: Across all three tasks, our goal is to manipulate an output image, shifting the bias from class $\mathbb{A}$ to $\mathbb{B}$ by utilising $\mathcal{S}$ variables to control manipulation severity. Our fundamental, precise prompt engineering results in Table 1 show that the relationships between $\mathcal{S}$ and our evaluation metrics are consistent for all four class combinations, evidencing that we have achieved fine control over the language model embedding output without re-training the model. We also see that extrapolating the embedding beyond $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ is effective as visualized in Fig. 5 and supported quantitatively in Table 1. For $\mathcal{S}<0$ and $\mathcal{S}>1$, it is evident that the embedding is still within clusters $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ respectively. Where $\mathcal{P}_{\mathbb{A}}\cong\mathcal{P}_{\mathbb{B}}\cong 0.5$, this implies that applying the relevant transformation with the corresponding $\mathcal{S}_{i}$ has shifted $\mathbf{X}$ to the border of clusters $\mathbb{A}$ and $\mathbb{B}$ i.e. an empirically-derived “half dog/half cat”.
@@
Through results of the $N_{20}$ and $N_{100}$ social bias tuning experiments in Table 2, we immediately observe that $\mathcal{S}_{i}$ values are dependent on the amount of seeds used for image generation. Covering a broader range of samples improves generalizability but at the cost of computation time. We opt for using labels to describe $\mathcal{S}_{i}$ and $\mathcal{P}_{i}$ for clarity in Table 2. However this is just a matter of syntax and $\mathcal{S}_{male}\rightarrow\mathcal{S}_{asian}$ are identical to $\mathcal{S}_{1}\rightarrow\mathcal{S}_{7}$ derived previously.
In Table 2, we report the bias mitigation results for each social marker (gender, age, race) and compare it to the base model, where $\mathcal{S}_{1\rightarrow 7}=0$. For the $N_{100}$ case, we see that gender biases were evenly distributed in the base model, without making any changes to $\mathcal{S}_{male}$, $\mathcal{S}_{female}$. Whereas for the $N_{20}$ case, there is a clear gender imbalance that required tuning. As predicted, changing $\mathcal{S}_{i}$ variables will have an impact on class representations at each tuning stage, leading to the final step where the biases are more evenly distributed. Related work by Clemmer et al. reported less than 2.5% gender bias and less than 10% ethnicity biases [10]. Feng et al. reported a convergence of around 50% gender probability distribution for binary gender classes (less than 7% bias) in [12]. Comparing our results in Table 2, we see that our mitigation method is comparable to the state of the art, as we report less than 10% gender, age and racial bias.
@@
Table 2: Social bias mitigation results. For $N_{20}$ and $N_{100}$ random seed experiments, we tune $\mathcal{S}_{i}$ to balance gender, age and racial representations, where $\mathcal{S}_{i}$ represents how much the embedding has been shifted toward the $i^{th}$ centroid based on Eq. (5). We exploit a CLIP VC model to report the frequency/probability ‘$\mathcal{P}_{i}$’ of generated images being classified as class $i$. Fig. 4 presents a visual representation $\mathcal{S}_{i}$ optimization.
Dog - Cat Bird - Plane Horse - Deer Car - Truck
$\mathbf{x}_{T}$ Photo Picture Image View - Photo Picture Image View - Photo Picture Image View - Photo Picture Image View
$\mathcal{S}$ 0 0.5 0.75 1 1.25 0 0.5 0.75 1 1.25 0 0.5 0.75 1 1.25 0 0.5 0.75 1 1.25
ASRVC 0 0.049 0.848 1 1 0.005 0.140 0.929 1 0.989 0 0 0.249 0.9 1 0 0.1 0.749 1 0.999
ASRVL 0.011 0.115 0.860 0.987 0.982 0.043 0.127 0.626 0.400 0.495 0 0.001 0.006 0.400 0.863 0.024 0.186 0.599 1 0.978
$\mathcal{P}_{\mathbb{B}}$ 0.000 0.059 0.837 0.988 0.993 0.005 0.141 0.921 0.999 0.980 0.000 0.003 0.244 0.916 0.996 0.001 0.119 0.731 0.980 0.993
@@
Table 3: Semantically-null backdoor attack results. For the four tasks, we define semantically-null trigger clusters $\mathbb{A}_{i}$, each hosting a unique severity value $\mathcal{S}_{i}$. Upon detection of a trigger ‘$x_{T_{i}}$’ in the prompt, we apply the shift to the embedding output based on Eq. (7). We report VL and VC attack success rate (ASRVL and ASRVC, respectively) and classifier confidence/probability for class $\mathbb{B}$ $\mathcal{P}_{\mathbb{B}}$, where $\mathcal{P}_{\mathbb{A}}=1-\mathcal{P}_{\mathbb{B}}$.
@@
Given the attack nature of backdoors, ASR metrics are important to demonstrate that our approach is effective. Logically, we hypothesized that ASRVL, ASRVC and $\mathcal{P}_{\mathbb{B}}$ are all proportional to $\mathcal{S}$ in the range $0\leq\mathcal{S}\leq 1$ and thus, $\mathcal{P}_{\mathbb{A}}~{}\frac{1}{\propto}~{}\mathcal{S}$ in the range $0\leq\mathcal{S}\leq 1$. In Table 3, we observe that our hypotheses were correct for all four class pairs. Through Fig. 6, as expected, the output images converge on their respective $c_{\mathbb{B}}$ as $\mathcal{S}\rightarrow 1$. The weakest performing task was the Horse-Deer attack, where we observe that unlike the others, it is the only case where ASRVC is not 100% at $\mathcal{S}=1$ This may be due to the similarities between the two labelled points. For others, we see that the attack is quite effective even at $\mathcal{S}=0.75$. Overall, we demonstrate an effective, computationally-efficient backdoor injection method, where we manipulate $\mathbf{X}$ and thus, shift T2I model biases based on the severity $\mathcal{S}_{i}$ and a semantically-null trigger space $\mathbb{A}_{i}$.
Figure 6: Qualitative backdoor attack results. We prepend $x_{T_{i}}$ to the base input prompt in our experiments e.g. “a photo of…” (a) $\mathbb{A}$ = horse, $\mathbb{B}$ = deer. (b) $\mathbb{A}$ = car, $\mathbb{B}$ = truck.
@@
For comparison, in [35], Vice et al. reported a backdoor ASR range of 47.2% to 87.9%. Chen et al. achieved a 79.3% to 99.6% ASR in [6] when targeting the diffusion/generative model component. Struppek et al. targeted embedded language models and reported an attack accuracy of around 90% [33]. Zhai et al. reported a 60.1% to 98.8% ASR range when conducting their backdoor implementations [38]. Thus, not only can we effectively control the severity of our attacks in real-time, achieving mean ASRVC and ASRVL scores of 97.5% and 69.2% respectively proves that at $\mathcal{S}=1.0$, our attacks are comparable to state-of-the-art works. Furthermore, by extrapolating $\mathbf{X}$ beyond $c_{\mathbb{B}}$ at $\mathcal{S}=1.25$, our reported ASRVC is higher than other methods [6, 33, 35, 38]. Finally, our backdoor attack does not require training or fine-tuning, unlike other methods [6, 33, 35, 38].
Limitations: Navigating the embedding space using labelled points as coordinates may cause a loss of image quality and production of generative artifacts based on the relationship of entities within the embedding space. For bias mitigation tasks, while demographic biases may change from point $\mathbb{A}\rightarrow\mathbb{B}$, the inter-cluster region may contain stereotype-encoding features that could effect evaluations. The frequency and relative position of these features could be captured using image captioning and VQA models to query generated content. A limitation of using centroids as backdoor attack targets is that text-alignment is $\frac{1}{\propto}~{}\mathcal{S}$, intentionally shifting the output toward a centroid $c_{i}\in\mathbb{E}^{n\times m}$.
@@ 5 Conclusion
We have presented a computationally efficient, bias manipulation method, leveraging high-dimensionality clusters and vector algebra to shift embedding outputs and ultimately, control bias characteristics in text-to-image models. We demonstrate that the embedding space can be exploited for bias manipulation and mitigation tasks, without requiring fine-tuning or gaining access to network weights and parameters. Our approach is multi-faceted and we highlight applications for: (i) precise prompt engineering, (ii) social bias mitigation and, (iii) using semantically-null triggers to inject malicious backdoors. We demonstrate that bias manipulations do not have to be computationally taxing and our results indicate that depending on the application, our method is comparable to both state-of-the-art bias mitigation and backdoor attack methods.
@@ 6 Acknowledgements
Dr. Jordan Vice is supported by the National Intelligence and Security Discovery Research Grants (NISDRG) project number 20100007, funded by the Australian Government. Dr. Naveed Akhtar is a recipient of the Australian Research Council Discovery Early Career Researcher Award (project number DE230101058) funded by the Australian Government. Professor Ajmal Mian is the recipient of an Australian Research Council Future Fellowship Award (project number FT210100268) funded by the Australian Government.